Write the verbal phrase as an equation or an inequality.
Difference of x and thrity is equal to fifty
Sum of x and eleven is greater than or equal to forty
Product of three and x is less then fourteen
Quotient of x and 3.5 is equal to nine

Found 2 solutions by checkley77, jpython:
Answer by checkley77(12844)

x-30=50
x+11>=40
3x<14
x/3.5=9

Answer by jpython(1)